# Time-Varying Gaussian Process Bandits with Unknown Prior

### v3

@@ 8 Conclusions
Within this work, we proposed the first algorithm with provable regret bound, capable of solving time-varying GP bandits problem under unknown prior. One limitation of our work is that the time-dependent part of the regret bound scales with $\sqrt{|U|}$ that is the square root of number of candidate priors. This might make the method perform poorly if the number of candidate priors is extreme. Finding a way to improve that dependence consistutes an interesting direction for future work.
@@
Let us consider some fixed $p\in\mathcal{\mathcal{U}}$. Let us define a new set of i.i.d. random variables $v_{i}$ for $i=1,\dots,T$, where each of them follows the same $R$-subgaussian distribution as the noise. Since $v_{i}$ is $R$-subgaussian, for any $0<n\leq T$ we have:
@@
which is a well-known result in the literature (e.g. see Corollary 5.5 in Lattimore & Szepesvári (2020)). Let us sample each $v_{i}$ for $i=1,\dots,T$ before the experiment starts and let them be a hidden state of the environment. Whenever for the current iteration $t$, we have that the set $S_{t-1}^{p}$ is expanded, we set the noise of that iteration as $\epsilon_{t}=v_{i}$, where $i=|S_{t}^{p}|$. Such a constructed noise must have the same distribution as the original noise. For each $i=1,\dots,T$, let us now define the event $E_{n}$ as:
$E_{n}:\quad\left|\sum_{i=1}^{n}v_{i}\right|>\sqrt{\xi_{n}n}.$
We now have:
$P\left(\exists_{t=1,\dots,T}\left|\sum_{n\in S_{t}^{p}}\epsilon_{n}\right|>\sqrt{\xi_{n}|S_{t}^{p}|}\right)=P\left(\bigcup_{n=1}^{|S_{T}^{p}|}E_{n}\right)\leq P\left(\bigcup_{n=1}^{T}E_{n}\right)$
$\leq\sum_{n=1}^{T}P\left(E_{n}\right)\leq 2\sum_{t=1}^{T}\exp\left(\frac{-\xi_{t}}{2R^{2}}\right)=\sum_{t=1}^{T}\frac{6\delta_{B}}{|\mathcal{\mathcal{U}}|\pi^{2}t^{2}}\leq\frac{\delta_{B}}{|\mathcal{\mathcal{U}}|},$
where the first inequality is due to the fact that $\bigcup_{n=1}^{|S_{T}^{p}|}E_{n}\subseteq\bigcup_{n=1}^{T}E_{n}$ as $|S_{T}^{p}|\leq T$. We finish the proof by taking a union bound over all $p\in\mathcal{\mathcal{U}}$. ∎
@@
$\sum_{t\notin\mathcal{C}}\beta_{t}^{+}\sigma^{+}_{t-1}(\bm{x}_{t})\leq\beta_{T}\sqrt{(T-|\mathcal{C}|)\sum_{t\notin\mathcal{C}}(\sigma^{+}_{t-1})^{2}(\bm{x}_{t})}\leq\beta_{T}\sqrt{T\sum_{p\in\mathcal{U}}\sum_{\begin{subarray}{c}t\notin\mathcal{C}\\ p_{t}=p\end{subarray}}(\sigma^{p}_{t-1})^{2}(\bm{x}_{t})},$ (5)
@@
$(\sigma^{u}_{A})^{2}(\bm{x})=k(\bm{x},\bm{x})^{u}-\bm{k}^{u}_{A}(\bm{x})^{T}(K^{u}_{A}+\sigma^{2}\mathcal{I})^{-1}\bm{k}^{u}_{A}(\bm{x}),$
where $K^{u^{*}}_{A}=[k^{u^{*}}(\bm{x},\bm{x}^{\prime})]_{\bm{x},\bm{x}^{\prime}\in A}$ and $\bm{k}^{u^{*}}_{A}=[k^{u^{*}}(\bm{x},\bm{x}^{\prime})]_{\bm{x}^{\prime}\in A}$ Notice that under such notation $\sigma^{u}_{t-1}=\sigma^{u}_{X_{t-1}}$, where $X_{t-1}=\{\bm{x}_{i}\}_{i=1}^{t-1}$. Let us define and $A^{p}_{t-1}=\{((\bm{x}_{i},i),y_{i}):1\leq i\leq t-1,p_{t}=p\}$ to be the set of points queried when $p_{t}=p$. For any two sets $S,S^{\prime}$ such that $S^{\prime}\subseteq S$ we have $\sigma^{p}_{S^{\prime}}(\bm{x})\leq\sigma^{p}_{S}(\bm{x})$ for all $\bm{x}\in\mathcal{X}$. As $A^{p}_{t-1}\subseteq\mathcal{D}_{t-1}$, we have:
$(\sigma^{+}_{t-1})^{2}(\bm{x}_{t})=(\sigma^{p}_{\mathcal{D}_{t-1}})^{2}(\bm{x}_{t})\leq(\sigma^{p}_{A^{p}_{t-1}})^{2}(\bm{x}_{t})\leq C\log(1+R^{-2}(\sigma^{p}_{A_{t-1}^{p}}(\bm{x}))^{2}),$
@@
$\beta_{T}\sqrt{T\sum_{p\in\mathcal{U}}\sum_{\begin{subarray}{c}t\notin\mathcal{C}\\ u_{t}=p\end{subarray}}(\sigma^{p}_{t-1})^{2}(\bm{x}_{t})}\leq\beta_{T}\sqrt{CT\sum_{p\in\mathcal{U}}\sum_{\begin{subarray}{c}t\notin\mathcal{C}\\ u_{t}=p\end{subarray}}\log(1+R^{-2}(\sigma^{p}_{A_{t-1}^{p}}(\bm{x}))^{2})}$
@@
Note that if event of Theorem 3.1 holds, then when $t=0$, we have that for all $\bm{x}\in\mathcal{X}$ and all $t^{\prime}\in[T]$:
$|f(\bm{x},t^{\prime})-\mu^{p^{\star}}(\bm{x},t)|=|f(\bm{x},t^{\prime})-\mu^{p^{\star}}_{0}(\bm{x},t^{\prime})|\leq\beta^{p^{\star}}_{1}\sigma^{p^{\star}}_{0}(\bm{x},t^{\prime})\leq\beta^{p^{\star}}_{1}k^{p^{\star}}((\bm{x},t^{\prime}),(\bm{x},t^{\prime}))\leq\beta^{p^{\star}}_{1}$
As such:
$\max_{t\in[T]}\max_{\bm{x}\in\mathcal{X}}|f(\bm{x},t^{\prime})|\leq\beta_{1}+\max_{t^{\prime}\in[T]}\max_{\bm{x}\in\mathcal{X}}|\mu^{p^{\star}}(\bm{x},t^{\prime})|,$
and defining $B=\beta_{1}+\max_{t^{\prime}\in[T]}\max_{\bm{x}\in\mathcal{X}}|\mu^{p^{\star}}(\bm{x},t^{\prime})|$ completes the proof.
@@
Note that if event of Theorem 3.3 holds, then when $t=0$, we have that for all $\bm{x}\in\mathcal{X}$ and $t^{\prime}\in[T]$:
$|f(\bm{x},t^{\prime})-\mu^{p^{\star}}([\bm{x}]_{1},t^{\prime})=|f(\bm{x},t^{\prime})-\mu^{p^{\star}}_{0}([\bm{x}]_{1},t^{\prime})\leq\beta_{1}\sigma^{p^{\star}}([\bm{x}]_{1},t^{\prime})+1\leq\beta_{1}k^{p^{\star}}([\bm{x}]_{1},[\bm{x}]_{1},t^{\prime})+1\leq\beta_{1}+1$
As such:
$\max_{t^{\prime}\in[T]}\sup_{\bm{x}\in\mathcal{X}}|f(\bm{x},t^{\prime})|\leq\beta_{1}+\max_{t^{\prime}\in[T]}\max_{\bm{x}\in D_{t}}|\mu(\bm{x},t^{\prime})|+1\leq\beta_{1}+\max_{t^{\prime}\in[T]}\sup_{\bm{x}\in\mathcal{X}}|\mu(\bm{x},t^{\prime})|+1,$
and defining $B=\beta_{1}+\max_{t^{\prime}\in[T]}\sup_{\bm{x}\in\mathcal{X}}|\mu(\bm{x},t^{\prime})|+1$ completes the proof.
@@ Appendix D Regret Balancing for Unknown Prior
4:     Select point to query $\bm{x}_{t}=\underset{\bm{x}\in\mathcal{X}}{\arg\max}\textrm{ UCB}_{t-1}^{p_{t}}(\bm{x})$
5:     Query the black-box $y_{t}=f(\bm{x}_{t},t)$
@@
11:        $\mathcal{U}_{t+1}=\left\{p\in\mathcal{U}_{t}:L_{t}(u)+\frac{1}{|S_{t}^{p}|}\sum_{\tau\in S_{t}^{p}}\beta_{\tau}^{p}\sigma^{p}_{\tau-1}(\bm{x}_{\tau})\geq\max_{p^{\prime}\in\mathcal{U}_{t}}L_{t}(p^{\prime})\right\}$

### v4

@@ 8 Conclusions
Within this work, we proposed the first algorithm with provable regret bound, capable of solving time-varying GP bandits problem under unknown prior. One limitation of our work is that the time-dependent part of the regret bound scales with $\sqrt{|U|}$ that is the square root of number of candidate priors. This might make the method perform poorly if the number of candidate priors is extreme. Another limitation is the aforementioned over-optimisim, which prevents the application of the proposed algorithms in certain settings. Finding a way to address those limitations consistutes an interesting direction for future work.
Acknowledgements
We thank anonymous reviewers who gave useful comments. Juliusz Ziomek was supported by the Oxford Ashton-Memorial Scholarship and EPSRC DTP grant EP/W524311/1. Masaki Adachi was supported by the Clarendon Fund, the Oxford Kobe Scholarship, the Watanabe Foundation, and Toyota Motor Corporation.
@@ Acknowledgements
We note that in the sum $\sum_{t\in S_{T}^{p}}\epsilon_{t}$, the set $S_{T}^{p}$ can depend on the noise values. However, note that the decision of which prior to choose at a given iteration $t$ (and thus whether or not $t\in S_{T}^{p}$) is made before observing the function value at that step and thus while $\{\epsilon_{t}\}_{t=1}^{\tau-1}$ can affect whether or not $\epsilon_{\tau}$ is included in $S_{T}^{p}$, they cannot affect its value. As such, bounding such a sum is equivalent to bounding a sum of i.i.d. variable with a history-dependent stopping rule that tells us when to stop putting more variables into the sum. We will do this by showing the inequality holds with high probability for any possible stopping time simultanously, so we can be sure that no matter what stopping time is chosen, the inequality will hold with high probability. Let $\{v_{i}\}_{i=1}^{\infty}$ be a sequence of i.i.d. random variables, each being $R$-subgaussian. We then have that:
$P\left(\exists_{t\in S_{T}^{p}}\left|\sum_{i=1}^{t}\epsilon_{i}\right|\leq\sqrt{\xi_{t}t}\right)$ $=P\left(\exists_{0\leq n\leq T}\exists_{t\in[n]}\left|\sum_{i=1}^{t}v_{i}\right|\leq\sqrt{\xi_{t}t},|S_{T}^{p}|=n\right)$
$\leq P\left(\exists_{0\leq n\leq T}\exists_{t\in[n]}\left|\sum_{i=1}^{t}v_{i}\right|\leq\sqrt{\xi_{t}t}\right)$
$=P\left(\exists_{0\leq t\leq T}\left|\sum_{i=1}^{t}v_{i}\right|\leq\sqrt{\xi_{t}t}\right)$
$\leq\sum_{t\in[T]}P\left(\left|\sum_{i=1}^{t}v_{i}\right|\leq\sqrt{\xi_{t}t}\right)$
$\leq 2\sum_{t\in[T]}\exp\left(-\frac{\xi_{t}}{2R^{2}}\right)$
$=2\sum_{t\in[T]}\frac{3\delta}{\pi^{2}t^{2}|\mathcal{U}|}\leq 2\sum_{t=1}^{\infty}\frac{3\delta}{\pi^{2}t^{2}|\mathcal{U}|}=\frac{\delta}{|\mathcal{U}|}.$
In the penultimate transition above we used the fact that since $v_{i}$ is $R$-subgaussian, for any $n>0$ we have:
@@
which is a well-known result in the literature (e.g. see Corollary 5.5 in Lattimore & Szepesvári (2020)). Taking union bound over priors proves the statement:
$P\left(\exists_{p\in\mathcal{U}}\exists_{t\in S_{T}^{p}}\left|\sum_{i=1}^{t}\epsilon_{i}\right|\leq\sqrt{\xi_{t}t}\right)\leq\sum_{p\in\mathcal{U}}\frac{\delta}{|\mathcal{U}|}=\delta.$
Note that the statement is proven for any prior $p\in\mathcal{U}$ and any $t\in S_{T}^{p}$. As such, the statement of the Lemma is a finite sample, high-probability bound and not an asymptotic statement. ∎
@@
$\sum_{t\notin\mathcal{C}}\beta_{t}^{+}\sigma^{+}_{t-1}(\mathbf{x}_{t})\leq\beta_{T}\sqrt{(T-|\mathcal{C}|)\sum_{t\notin\mathcal{C}}(\sigma^{+}_{t-1})^{2}(\mathbf{x}_{t})}\leq\beta_{T}\sqrt{T\sum_{p\in\mathcal{U}}\sum_{\begin{subarray}{c}t\notin\mathcal{C}\\ p_{t}=p\end{subarray}}(\sigma^{p}_{t-1})^{2}(\mathbf{x}_{t})},$ (5)
@@
$(\sigma^{u}_{A})^{2}(\mathbf{x})=k(\mathbf{x},\mathbf{x})^{u}-\mathbf{k}^{u}_{A}(\mathbf{x})^{T}(K^{u}_{A}+\sigma^{2}\mathcal{I})^{-1}\mathbf{k}^{u}_{A}(\mathbf{x}),$
where $K^{u^{*}}_{A}=[k^{u^{*}}(\mathbf{x},\mathbf{x}^{\prime})]_{\mathbf{x},\mathbf{x}^{\prime}\in A}$ and $\mathbf{k}^{u^{*}}_{A}=[k^{u^{*}}(\mathbf{x},\mathbf{x}^{\prime})]_{\mathbf{x}^{\prime}\in A}$ Notice that under such notation $\sigma^{u}_{t-1}=\sigma^{u}_{X_{t-1}}$, where $X_{t-1}=\{\mathbf{x}_{i}\}_{i=1}^{t-1}$. Let us define and $A^{p}_{t-1}=\{((\mathbf{x}_{i},i),y_{i}):1\leq i\leq t-1,p_{t}=p\}$ to be the set of points queried when $p_{t}=p$. For any two sets $S,S^{\prime}$ such that $S^{\prime}\subseteq S$ we have $\sigma^{p}_{S^{\prime}}(\mathbf{x})\leq\sigma^{p}_{S}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}$. As $A^{p}_{t-1}\subseteq\mathcal{D}_{t-1}$, we have:
$(\sigma^{+}_{t-1})^{2}(\mathbf{x}_{t})=(\sigma^{p}_{\mathcal{D}_{t-1}})^{2}(\mathbf{x}_{t})\leq(\sigma^{p}_{A^{p}_{t-1}})^{2}(\mathbf{x}_{t})\leq C\log(1+R^{-2}(\sigma^{p}_{A_{t-1}^{p}}(\mathbf{x}))^{2}),$
@@
$\beta_{T}\sqrt{T\sum_{p\in\mathcal{U}}\sum_{\begin{subarray}{c}t\notin\mathcal{C}\\ u_{t}=p\end{subarray}}(\sigma^{p}_{t-1})^{2}(\mathbf{x}_{t})}\leq\beta_{T}\sqrt{CT\sum_{p\in\mathcal{U}}\sum_{\begin{subarray}{c}t\notin\mathcal{C}\\ u_{t}=p\end{subarray}}\log(1+R^{-2}(\sigma^{p}_{A_{t-1}^{p}}(\mathbf{x}))^{2})}$
@@
Using the same proof idea as Lemma 5.1. of Srinivas et al. (2010), where we simply replace the posterior mean $\mu_{t-1}$ and variance $\sigma^{2}_{t-1}$ with the prior mean $\mu$ and variance $\sigma$ we get that with probability at least $1-\delta_{c}$, for all $\mathbf{x}\in\mathcal{X}$ and all $t^{\prime}\in[T]$:
$|f(\mathbf{x},t^{\prime})-\mu^{p^{\star}}(\mathbf{x},t^{\prime})|\leq\beta^{p^{\star}}_{t}\sigma^{p^{\star}}_{0}(\mathbf{x},t^{\prime})\leq\beta^{p^{\star}}_{t}k^{p^{\star}}((\mathbf{x},t^{\prime}),(\mathbf{x},t^{\prime}))\leq\hat{\beta}_{t}^{p^{\star}},$
where $\beta^{p}_{t}=\sqrt{2\log\left(\frac{|\mathcal{X}|\pi^{2}t^{2}}{2\delta_{c}}\right)}$ As such:
$\max_{t^{\prime}\in[t]}\max_{\mathbf{x}\in\mathcal{X}}|f(\mathbf{x},t^{\prime})|\leq\beta^{p^{\star}}_{t}+\max_{t^{\prime}\in[t]}\max_{\mathbf{x}\in\mathcal{X}}|\mu^{p^{\star}}(\mathbf{x},t^{\prime})|\leq B_{t}.$
@@
Using the same proof idea as Appendix C.1. of Bogunovic et al. (2016), where we simply replace the posterior mean $\mu_{t-1}$ and variance $\sigma^{2}_{t-1}$ with the prior mean $\mu$ and variance $\sigma$ we get that with probability at least $1-\delta_{c}$, for all $\mathbf{x}\in\mathcal{X}$ and all $t^{\prime}\in[T]$:
$|f(\mathbf{x},t^{\prime})-\mu^{p^{\star}}([\mathbf{x}]_{1},t^{\prime})\leq\beta^{p^{\star}}_{t^{\prime}}\sigma^{p^{\star}}([\mathbf{x}]_{1},t^{\prime})+1\leq\beta^{p^{\star}}_{t^{\prime}}k^{p^{\star}}([\mathbf{x}]_{1},[\mathbf{x}]_{1},t^{\prime})+1\leq\beta^{p^{\star}}_{t^{\prime}}+1,$
where $\beta_{t}^{p}=\sqrt{2\log\left(\frac{3\pi^{2}t^{2}}{2\delta_{C}}\right)+2d\log\left(dt^{2}rb^{p}\sqrt{\log\left(\frac{3da^{p}}{2\delta_{C}}\right)}\right)}$. As such:
$\max_{t^{\prime}\in[t]}\sup_{\mathbf{x}\in\mathcal{X}}|f(\mathbf{x},t^{\prime})|\leq\beta^{p^{\star}}_{t^{\prime}}+\max_{t^{\prime}\in[t]}\max_{\mathbf{x}\in D_{t}}|\mu^{p^{\star}}(\mathbf{x},t^{\prime})|+1\leq\beta^{p^{\star}}_{t^{\prime}}+\max_{t^{\prime}\in[t]}\sup_{\mathbf{x}\in\mathcal{X}}|\mu^{p^{\star}}(\mathbf{x},t^{\prime})|+1\leq B_{t}.$
@@ Appendix D Regret Balancing for Unknown Prior
4:     Select point to query $\mathbf{x}_{t}=\underset{\mathbf{x}\in\mathcal{X}}{\arg\max}\textrm{ UCB}_{t-1}^{p_{t}}(\mathbf{x})$
5:     Query the black-box $y_{t}=f(\mathbf{x}_{t},t)$
@@
11:        $\mathcal{U}_{t+1}=\left\{p\in\mathcal{U}_{t}:L_{t}(u)+\frac{1}{|S_{t}^{p}|}\sum_{\tau\in S_{t}^{p}}\beta_{\tau}^{p}\sigma^{p}_{\tau-1}(\mathbf{x}_{\tau})\geq\max_{p^{\prime}\in\mathcal{U}_{t}}L_{t}(p^{\prime})\right\}$
@@
Appendix E Proof of Proposition 6.2
Using the same reasoning as in Appendix E of Bogunovic et al. (2016) we have that in squared exponential case $\gamma_{T}=\tilde{\mathcal{O}}(\max\{T\varepsilon^{1/3},1\})$ and in $\nu$-Matern case $\gamma_{T}=\tilde{\mathcal{O}}(\max\{T\varepsilon^{\frac{1-c}{3-c}},T^{c}\})$ with $c=\frac{d(d+1)}{2\nu+d(d+1)}$.
Appendix F Proof of Lemma 6.1
To prove the first statement we need to show that with probability 1, we have $\forall_{\epsilon>0}\exists_{T^{\prime}(\epsilon)>0}\forall_{T>T^{\prime}(\epsilon)}\frac{t^{\star}}{T^{\prime}(\epsilon)}<\epsilon$. Note that since confidence parameters $\xi_{t}$ and $\beta_{t}$ do not require the knowledge of $T$ beforehand, we can always just decide to continue run the algorithm for more iterations after exceeding some inital iteration number. Let $\tau_{i}$ be the iteration, where we reject the $i$th prior if we run the algorithm for at least $\tau_{i}$ iterations. Note that $(\tau_{1},\dots,\tau_{|\mathcal{U}|})$ are random variables that depend on the randomness in sampling black-box $f$ from the prior (and its random evolution through time), and on the randomness in noise $\{\epsilon_{t}\}_{t=1}^{\infty}$, but do not depend on any randomness in the algorithm. Thus conditioning on noise variables and black-box function and its evolution, we have that $(\tau_{1},\dots,\tau_{|\mathcal{U}|})$ are determinstic variables. Let us denote $\tau_{i}=\infty$ if less than $i$ priors were rejected. Thus $t^{\star}\leq\max_{i=1,\dots,|\mathcal{U}|}\{\tau_{i}|\tau_{i}\neq\infty\}$ and conditioned on the aforementioned randomness $t^{\star}$ also becomes deterministic. As such, setting $T^{\prime}(\epsilon)=\frac{2\max_{i=1,\dots,|\mathcal{U}|}\{\tau_{i}|\tau_{i}\neq\infty\}}{\epsilon}$, we get that:
$P\left(\forall_{T>T^{\prime}(\epsilon)}\frac{t^{\star}}{T}\leq\frac{\max_{i=1,\dots,|\mathcal{U}|}\{\tau_{i}|\tau_{i}\neq\infty\}}{T}<\epsilon\Big{|}\{f(x_{t},t)\}_{t=1}^{\infty},\{\epsilon_{t}\}_{t=1}^{\infty}\right)=1,$
and as this construction can be done for any $\epsilon>0$, we have:
$P\left(\forall_{\epsilon>0}\forall_{T>T^{\prime}(\epsilon)}\frac{t^{\star}}{T}<\epsilon\Big{|}\{f(x_{t},t)\}_{t=1}^{\infty},\{\epsilon_{t}\}_{t=1}^{\infty}\right)=1.$
[…]
$P\left(\forall_{\epsilon>0}\forall_{T>T^{\prime}(\epsilon)}\frac{t^{\star}}{T}<\epsilon\right)=1.$
This finishes the proof for the first statement. To prove the second statement, we can use the same reasoning as above and set $T^{\prime}(\epsilon)=\frac{2B_{t^{\star}}}{\epsilon}$, as since $t^{\star}$ is deterministic, so is $B_{t^{\star}}$
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes (regret analysis)]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]